OpenStudy (anonymous):

You have 3/4 of a pie leftover after your party. Each guest takes home 1/6 of the leftover pie. What fraction of the pie does each guest take home?

OpenStudy (one098):

Okay, first make each fraction have the same denominator. Can you do that?

OpenStudy (one098):

\[\frac{ 3 }{ 4 }\times \frac{ 3 }{ 3 }\] \[\frac{ 1 }{ 6 }\times \frac{ 2 }{ 2 }\]

OpenStudy (one098):

Once you do that, both fractions will have the same denominator.

OpenStudy (anonymous):

so 9/12 and 2/12

OpenStudy (one098):

Okay, right.

OpenStudy (one098):

So now, subtract 9-2.

OpenStudy (anonymous):

7

OpenStudy (one098):

>_< Sorry, what you have to do is multiply both fractions. \[\frac{ 3 }{ 4 }\times \frac{ 1 }{ 6 }\]What did you get?

OpenStudy (anonymous):

3/24

OpenStudy (one098):

Now simplify.

OpenStudy (anonymous):

can you even simplify that

OpenStudy (one098):

Yes, of course. Can 3 go into 24?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

8 times

OpenStudy (one098):

Exactly, so if you do: \[\frac{ 3 }{ 24 }\div \frac{ 3 }{ 3 }=?\]

OpenStudy (anonymous):

1/8

OpenStudy (one098):

Right! So that's your answer.

OpenStudy (anonymous):

ok thanks

OpenStudy (one098):

You're welcome!